# Asymptotically conformal CFL quark matter within a nonlocal chiral quark model

### v1

@@ III Nonlocal NJL model at high densities
Thus, the vector field asymptotic is divergent in the local case. This is in a striking disagreement with the nonlocal NJL model providing a convergent $\omega_{\infty}$. Furthermore, in the local case the quark contribution to the thermodynamic potential scales as $(\mu+\omega_{\infty})^{4}\propto\mu^{4/3}$ and is negligible compared to the contribution of the vector field $-\omega^{2}/4G_{V}\propto\mu^{2}$. This leads to the pressure $p=\mu^{2}/4G_{V}$, speed of sound $c_{S}^{2}=1$ and dimensionless interaction measure $\delta=-2/3$ violating the conformal behavior.
It is also interesting to consider another version of the NJL model with the scalar and vector interactions treated nonlocally but the diquark pairing one kept local. In this case the formfactor in Eq. (25) should be set $g_{\bf k}=1$ and the effects of vector and scalar interactions can be neglected since $\omega_{\infty}$ and $\sigma_{\infty}$ are finite and vanish, respectively. As it is shown below, the pairing gap diverges if the corresponding interaction is treated locally. Moreover, $\Delta_{\infty}\gg\mu$ in this case, which allows us to approximate $\epsilon^{\pm}_{j{\bf k}}\simeq\mp\Delta\zeta_{j}$. This cancels the zero point terms of quarks and antiquarks in Eq. (25) and, unlike the nonlocal NJL model, yields the divergent $\Delta_{\infty}=G_{D}\tilde{d}\mu^{3}/3\pi^{2}$. As expected, quark chemical potential is negligibly small compared to this asymptotic of the quark pairing gap. In this case the quark contribution to the thermodynamic potential reads $-\tilde{d}\Delta_{\infty}\mu^{3}/6\pi^{2}$, while the resulting pressure, speed of sound and dimensionless interaction measure attain the nonconformal values $p=G_{D}\tilde{d}^{2}\mu^{6}/36\pi^{4}$, $c_{S}^{2}=1/5$ and $\delta=2/15$.
@@ IV Neutron stars with CFLL quark cores
Quark deconfinement significantly impacts the phenomenology of NSs. It can be triggered by a mechanical instability of the NS matter against the gravitational compression leading to the responseless density pile up followed by the Mott dissociation of hadrons. Recently BEC condensation of sexaquarks was proposed as a mechanism of such instability Blaschke et al. (2022); Shahrbaf et al. (2022). Indeed, under the color gauge group diquarks transform as antitriples, thus, making the QCD interaction between them analogous to the one between antiquarks. This potentially allows formation and BEC of a deeply bound sexaquark state of three CFL diquarks qualitatively different from the dihyperon state of the same flavor content Jaffe (1977). Unlike the dihyperon state, the sexaquark is not yet studied in lQCD. This makes us to rely on the available results obtained with QCD sum rules Azizi et al. (2020) and a constituent quark model Buccella (2020), which predict $m_{S}=1180^{+40}_{-28}$ MeV and $m_{S}=1880\pm 3$ MeV, respectively. Despite significantly differing, these results support strong and weak stability of sexaquarks. Following Refs. Blaschke et al. (2022); Shahrbaf et al. (2022), we prefer a conservative estimate of the sexaquark mass as the upper limit of the threshold with respect to the dominant weak process being the decay to electron, proton and $\Lambda$-baryion, i.e. $m_{S}=m_{e}+m_{p}+m_{\Lambda}=2054$ MeV. Such sexaquark can decay only in a doubly-weak channel, which makes its lifetime compatible to the age of the Universe Farrar and Zaharijas (2004), so providing the effective stability on the timescales corresponding to NSs. For the sake of simplicity we also ignore medium modification of the sexaquark mass caused by the Pauli blocking among the constituent quarks. Accounting for such a dependence, however, would not drastically change the considered scenario. It only leads to a moderate density dependence of $m_{S}$ and insignificantly increases ability of the sexaquark enriched NS matter to resist the gravitational compression Shahrbaf et al. (2022).
@@
The flavor content of sexaquarks $uuddss$ coincides with the one of three diquarks $ud$, $ds$ and $su$ being a Bardeen–Cooper–Schrieffer (BCS) type quark correlations. Thus, dissociation of sexaquarks causes formation of a color-superconducting three-flavor state of quark matter dominated by the above diquark correlations, i.e. the CFLL quark matter Blaschke et al. (2022). This state exists in full thermodynamic equilibrium with the nuclear matter, i.e. their pressures coincide, while the common chemical potential $\mu_{B}=m_{S}/2$ is defined by the condition of the BEC of sexaquarks Blaschke et al. (2022). This is formally equivalent to the Maxwell construction of the deconfinement phase transition. The main difference is that within the considered scenario the position of deconfinement is unambiguously defined by the sexaquark mass. For a given EoS of nuclear matter and a fixed vector coupling this requires adjusting the diquark one. Thus, below we treat $\eta_{V}$ as a single independent parameter of the model. It is known that increasing the vector coupling increases the onset density of quark matter, while the diquark one causes the opposite effect (see the corresponding discussion e.g. in Refs. Baym et al. (2018); Ivanytskyi and Blaschke (2022a, b)). Therefore, having a fixed value of the chemical potential of quark deconfinement defined by the sexaquark mass requires $\eta_{D}$ to be a growing function of $\eta_{V}$. This behavior is seen from Table 1, which summarizes the hybrid EoS used below. Since the considered scenario assumes deconfinement of a color-superconducting CFLL quark matter, the diquark coupling obtained from the requirement that quark-hadron transition onsets at the BEC of sexaquarks should provide a finite diquark pairing gap found as a solution of Eq. (25) at $\mu_{B}=m_{S}/2$. Within the considered scenario transition from hadron to quark matter in the general case is of the first order. It is indicated by a finite jump of the baryon density. Requiring this density jump across the transition from nuclear to quark matter to be non-negative, we limit the ranges of the vector and diquark couplings from above. The vector and diquark couplings producing the EoSs summarized in Table 1 respect these two requirements. They also assure the absence of the vacuum color superconductivity since they are always below the critical value $\eta_{D}=0.765$.
$\eta_{V}$ $\eta_{D}$ $\rm\varepsilon_{\rm max}~{}[MeV/fm^{3}]$ $\rm M_{\rm max}~{}[M_{\odot}]$ $\rm R_{1.4}$ [km] $\Lambda_{1.4}$
0.55 0.27 1281 1.94 12.23 280
0.70 0.34 1203 2.01 12.39 321
0.85 0.38 1164 2.09 12.58 393
1.00 0.41 1086 2.15 12.80 424
1.21 0.44 1038 2.24 13.15 569
Table 1: Vector $\eta_{V}$ and diquark $\eta_{D}$ couplings producing the EoSs shown in Fig. 1, maximal energy densities $\varepsilon_{\rm max}$ reached in the centers of the heaviest NSs, as well as the corresponding maximum masses $\rm M_{\rm max}$, radii $\rm R_{1.4}$ and tidal deformabilities $\Lambda_{1.4}$ of a canonical NS extracted from Figs. 3 and 4.
In order to model NSs we construct a family of hybrid quark-hadron EoSs based on the procedure described above. The hadron part is modelled by the DD2npY-T EoS supplemented with a crust EoS Typel et al. (2010); Shahrbaf et al. (2022). It includes nucleons and hyperons, is consistent with the low density constraint of the chiral effective field theory (CEFT) Krüger et al. (2013) and reproduces the main properties of the nuclear matter ground state. Using the sexaquark mass $m_{S}=2054$ MeV the DD2npY-T EoS yields the quark onset density $n_{B}=0.25~{}{\rm fm}^{-3}$ and the corresponding energy density $\varepsilon=245~{}{\rm MeV~{}fm}^{-3}$. The resulting EoSs in the plane of pressure versus energy density are shown in Fig. 1. It agrees well with the constraints extracted from the observational data of the pulsars PSR J1614+2230 Hebeler et al. (2013) and PSR J0740+6620 Miller et al. (2021). An insignificant tension with the PSR J0740+6620 constraint is due to the absence of a first order phase transition in the analysis of Ref. Miller et al. (2021), while the corresponding observational data are fully reproduced within the present model. The flat parts of the depicted curves correspond to the mixed quark-hadron phase where the BEC of sexaquarks exists. As is seen, increasing the vector coupling stiffens the quark branch of the EoS and simultaneously reduces the density jump across the quark-hadron transition. The maximal considered value of $\eta_{V}=1.21$ corresponds to vanishing of this jump. At sufficiently large values of this parameter the quark branch of the hybrid EoS is stiffer than the DD2npY-T EoS at the energy densities above $400~{}{\rm MeV~{}fm}^{-3}$. As is shown below, this allows the hybrid EoSs to reach the maximum NS masses above the value $2.1~{}{\rm M}_{\odot}$ corresponding to the DD2npY-T EoS. It is also seen from Fig. 1 that the developed EoSs are consistent with the constraint of perturbative QCD Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018). This is provided by the nonlocal treatment of quark interactions allowing the present model to asymptotically reach the conformal limit being a subject for QCD.
@@
Figure 1: Hybrid EoS of cold electrically neutral quark-hadron matter at $\beta$-equilibrium in the plane of pressure $p$ versus energy density $\varepsilon$. The empty white circle indicates onset of hyperons in pure hadron matter. The calculations are performed for the sexaquark mass $m_{S}=2054$ MeV. The nuclear matter Krüger et al. (2013), astrophysical Miller et al. (2021); Hebeler et al. (2013) and perturbative QCD Fraga et al. (2014) constraints represented by the shaded areas are discussed in the text.
@@
The speed of sound and dimensionless interaction measure of the present model are shown in Fig. 2. The hadron branch of the hybrid EoS demonstrates monotonously growing $c_{S}^{2}$ and decreasing $\delta$. In the purely hadronic case $c_{S}^{2}$ is nonmonotonous due to the onset of hyperons. In the mixed phase region the speed of sound vanishes due to the constant pressure while the dimensionless interaction measure exhibits a non-smooth behavior caused by a discontinuous density jump. Note, at the limiting value of the vector coupling $\eta_{V}=1.21$ the first order quark-hadron transition degenerates to the second order one indicated by the vanishing density jump. In this case the region with vanishing $c_{S}^{2}$ disappears, while preserving a discontinuity of this quantity. At this value of the vector coupling $\delta$ is a smooth function of $\varepsilon$. The quark branch of the hybrid EoS exhibits a nonmonotonous behavior of the speed of sound and dimensionless interaction measure. Right after the quark deconfinement $c_{S}^{2}$ increases and $\delta$ decreases with the energy density growth. This signals stiffening of the EoS caused by the quark deconfinement. The larger the vector coupling, the stiffer is the quark branch of the EoS. This is manifested by increasing $c_{S}^{2}$ and decreasing $\delta$ at growing $\eta_{V}$. The speed of sound develops a peak at $\varepsilon\simeq 500~{}{\rm MeV/fm}^{3}$. Thus, $c_{S}^{2}$ reaches its maximal value in the quark matter phase,
@@
Precursors of such a behavior of $c_{S}^{2}$ and $\delta$ in the centers of the heaviest NSs were discussed within the context of approximately restored conformal symmetry Fujimoto et al. (2022); Annala et al. (2023). At the same time, the conformality of a theory assumes that the corresponding action is scale invariant. This implies an absence of dimensionfull parameters in the action and leads to $p\propto\varepsilon\propto\mu^{4}$ just due to the dimensionality reason. In this case $c_{S}^{2}=1/3$ and $\delta=0$ follow as a consequence of the scale invariance. However, the inverse statement that $c_{S}^{2}=1/3$ and $\delta=0$ at a certain density imply the scale invariance is not correct in the general case. A simple counterexample is given by the model of ultraheavy quarks with local vector repulsion and a constant negative bag-like pressure shift. This model obviously breaks the scale invariance by three dimensionful parameters, i.e. by the quark mass, the vector coupling and the bag constant. The interested readers can find the corresponding consideration in Appendix A. At the same time, at the non-perturbative regimes the CFLL quark matter is characterized by the effective quark mass, vector field amplitude and pairing gap. These dynamically generated parameters are not negligible compared to the quark chemical potential and break the scale invariance of the effective action. For example, at $\eta_{V}=1$ and $\varepsilon=2100~{}\rm MeV/fm^{3}$ one gets $\omega/\mu=0.58$, while the speed of sound and dimensionless interaction measure are close to the conformal values. Therefore, the question about an approximately conformal behavior of quark matter at non-perturbative densities can not be addressed based just on the analysis of its EoS and requires an independent consideration of the effective action and relevant microscopic quantities in the context of their (im)possible scale invariance. Such a study is beyond the scope of the present paper and should be performed separately.
@@
Fig. 2 also demonstrates that the scenario of the CFLL quark matter in NS agrees with the constraints on the speed of sound and dimensionless interaction measure of the NS matter. These constraints were extracted from the astrophysical data based on the model agnostic statistical analysis at the 94 % confidence level Altiparmak et al. ; Marczenko et al. . The mentioned agreement is rather remarkable since unlike the present work the analysis from Refs. Altiparmak et al. ; Marczenko et al. does not explicitly include the possibility of a first order phase transition. A possible phenomenological explanation may be related to an early quark deconfinement at an onset mass below the masses of the analyzed NSs. In this case all these astrophysical objects lay on the quark branch of the mass-radius relation and, thus, are weakly sensitive to the details of the deconfinement phase transition. It is also seen from Fig. 2 that the speed of sound of the nonlocal NJL model is consistent with the perturbative QCD predictions Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018) at the energy densities above $\rm 10^{5}~{}MeV/fm^{3}$. In the case of the dimensionless interaction measure the agreement requires even larger $\varepsilon$. As discussed in Section III, this is due to the impossibility to adjust the formfactor according to the perturbative QCD asymptotics of the speed of $c_{S}^{2}$ and $\delta$. A better agreement can be reached by the piecewise $g_{\bf k}$ being Gaussian at small momenta and behaving as $g_{\bf k}\propto\alpha_{s}/{\bf k}^{2}$ at large momenta Hell et al. (2011); Kashiwa et al. (2011).
@@
As was shown in Section II, the effective quark mass, vector field amplitude and pairing gap are negligible at high densities. This insures the conformal behavior of the CFLL quark matter signalled by the speed of sound and dimensionless interaction measure assymptoticlly converging to the conformal values $c_{S}^{2}=1/3$ and $\delta=0$. Fig. 2 demonstrates this. It is also seen that convergence to the conformal limit is delayed by increasing the vector coupling. This might look paradoxical since for the Gaussian formfactor and a finite current quark mass the high density behavior of $c_{S}^{2}\simeq 1/3-m^{2}/3\mu^{2}$ and $\delta\simeq m^{2}/3\mu^{2}$ is controlled by $m$ only and so all the curves in Fig. 2 should converge to a universal one independent on $\eta_{V}$. However, the above relations are valid only when the vector repulsion term in Eq. (34) is negligible compared to the mass one. For the present choice of the formfactor this requires the quark chemical potential to fulfill the requirement $\mu^{3}g_{\mu}\ll-m^{2}\Lambda^{2}/\omega_{\infty}$. It is clear that this condition is strongly dependent on the vector coupling appearing through the asymptotic value of the vector field. For example, at $\eta_{V}=0.55$ and $\eta_{V}=1.00$ this requires $\mu_{B}\gg 6439$ MeV, $\varepsilon\gg 5.10\cdot 10^{7}~{}\rm MeV/fm^{3}$ and $\mu_{B}\gg 6586$ MeV, $\varepsilon\gg 5.58\cdot 10^{7}~{}{\rm MeV/fm}^{3}$, respectively.
@@
Comparing the speed of sound and dimensionless interaction measure of the present model and the NJL model with local vector repulsion from Ref. Blaschke et al. (2022) allows us to analyze the phenomenological influence of the nonlocal treatment of the mentioned interaction channel. Fig. 2 shows $c_{S}^{2}$ and $\delta$ of pure quark matter obtained for the model with the local vector interaction. Note, in this case significant stiffness of the quark matter EoS excludes the possibility of its Maxwell crossing with the hadron EoS used in this paper  [^2] The review Baym et al. (2018) provides an excellent discussion of how the stiffness of a quark EoS interplays with the Maxwell construction of the quark-hadron phase transition. The first order phase transition from chirally broken normal quark matter to partially chirally symmetric color-superconducting quark matter is indicated in Fig. 2 by discontinuous jumps of $c_{S}^{2}$ and kinks of $\delta$. The stiffness of the quark matter EoS mentioned above leads to large values of the speed of sound always exceeding the one corresponding to the nonlocal NJL model. As is seen, unlike the nonlocal NJL model, $c_{S}^{2}$ monotonously grows asymptotically reaching the speed of light (see the discussion in Section III). As argued above, the local treatment of the vector repulsion does not allow the NJL model to reach the conformal limit. The same conclusion can be drawn from the behavior of the dimensionless interaction measure. At the local vector repulsion it monotonously decreases saturating to the nonconformal value $-2/3$. It is also worth mentioning that in the case of the local treatment of the vector repulsion the quark matter EoS does not exhibit a region where $c_{S}^{2}$ and $\delta$ simultaneously attain nearly conformal values.

### v2

@@ III Nonlocal NJL model at high densities
Thus, the vector field asymptotically diverges in the local case. This is in a striking disagreement with the nonlocal NJL model providing a convergent $\omega_{\infty}$. Furthermore, in the local case the quark contribution to the thermodynamic potential scales as $(\mu+\omega_{\infty})^{4}\propto\mu^{4/3}$ and is negligible compared to the contribution of the vector field $-\omega^{2}/4G_{V}\propto\mu^{2}$. This leads to the pressure $p=\mu^{2}/4G_{V}$, speed of sound $c_{S}^{2}=1$ and dimensionless interaction measure $\delta=-2/3$ violating the conformal behavior.
It is also interesting to consider another version of the NJL model with the scalar and vector interactions treated nonlocally but the diquark pairing kept local. In this case the formfactor in Eq. (25) should be set $g_{\bf k}=1$ and the effects of vector and scalar interactions can be neglected since $\omega_{\infty}$ is finite and $\sigma_{\infty}$ vanishes. As it is shown below, the pairing gap diverges if the corresponding interaction is treated locally. Moreover, $\Delta_{\infty}\gg\mu$ in this case, which allows us to approximate $\epsilon^{\pm}_{j{\bf k}}\simeq\mp\Delta\zeta_{j}$. This cancels the zero point terms of quarks and antiquarks in Eq. (25) and, unlike the nonlocal NJL model, yields the divergent $\Delta_{\infty}=G_{D}\tilde{d}\mu^{3}/3\pi^{2}$. As expected, quark chemical potential is negligibly small compared to this asymptotote of the quark pairing gap. In this case the quark contribution to the thermodynamic potential reads $-\tilde{d}\Delta_{\infty}\mu^{3}/6\pi^{2}$, while the resulting pressure, speed of sound and dimensionless interaction measure attain the nonconformal values $p=G_{D}\tilde{d}^{2}\mu^{6}/36\pi^{4}$, $c_{S}^{2}=1/5$ and $\delta=2/15$.
@@ IV Neutron stars with CFLL quark cores
Quark deconfinement significantly impacts the phenomenology of NSs. The most common way of considering this scenario corresponds to constructing a two-phase EoS of the NS matter by matching within the Maxwell construction independent quark and hadron EoSs. This scheme does not address the microscopic mechanism of conversion of hadrons to quarks. While this mechanism lays beyond the scope of the present paper, we consider the recently proposed picture when BEC condensation of sexaquarks causes a mechanical instability of the NS matter against the gravitational compression and leads to a responseless density pile up followed by the Mott dissociation of hadrons and deconfinement of quarks Blaschke et al. (2022); Shahrbaf et al. (2022). It is worth mentioning that the nonlocal NJL model developed in Section III is not limited to this picture, which is considered for the sake of illustration of the possible mechanism of quark deconfinement in NSs.
@@
Under the color gauge group diquarks transform as antitriples, thus, making the QCD interaction between them analogous to the one between antiquarks. This potentially allows formation and BEC of a deeply bound sexaquark state of three CFL diquarks qualitatively different from the dihyperon state of the same flavor content Jaffe (1977). Unlike the dihyperon state, the sexaquark is not yet studied in lQCD. Therefore we rely on the results obtained with QCD sum rules Azizi et al. (2020) and a constituent quark model Buccella (2020), which predict $m_{S}=1180^{+40}_{-28}$ MeV and $m_{S}=1880\pm 3$ MeV, respectively. Despite being significantly different, these results support stability of sexaquarks with respect to its strong and weak decays. Following Refs. Blaschke et al. (2022); Shahrbaf et al. (2022), we prefer a conservative estimate of the sexaquark mass as the upper limit of the threshold with respect to the dominant weak process being the decay to electron, proton and $\Lambda$-baryion, i.e. $m_{S}=m_{e}+m_{p}+m_{\Lambda}=2054$ MeV. Such sexaquark can decay only in a doubly-weak channel, which makes its lifetime compatible to the age of the Universe Farrar and Zaharijas (2004), so providing the effective stability on the timescales corresponding to NSs. For the sake of simplicity we also ignore medium modification of the sexaquark mass caused by the Pauli blocking among the constituent quarks. Accounting for such a dependence, however, would not drastically change the considered scenario. It only leads to a moderate density dependence of $m_{S}$ and insignificantly increases ability of the sexaquark enriched NS matter to resist the gravitational compression Shahrbaf et al. (2022).
The described sexaquark is included to the hadron part of the hybrid quark-hadron EoS used below within the DD2npY-T model supplemented with a crust EoS Shahrbaf et al. (2022). It also includes nucleons and hyperons, is consistent with the low density constraint of the chiral effective field theory (CEFT) Krüger et al. (2013) and reproduces the main properties of the nuclear matter ground state. At $m_{S}=2054$ MeV the BEC of sexaquarks onsets at the baryon density $n_{B}=0.25~{}{\rm fm}^{-3}$ and the energy density $\varepsilon=245~{}{\rm MeV~{}fm}^{-3}$. The corresponding chemiocal potential $\mu_{B}=m_{S}/2$ is defined by the condition of the BEC of sexaquarks Blaschke et al. (2022); Shahrbaf et al. (2022).
Further increase of density of the hadron matter occurs due to populating the BEC of sexaquarks, which leaves the pressure and chemical potential unchanged Shahrbaf et al. (2022). Thus, the BEC of sexaquarks causes mechanical instability of hadron matter being compressed in the gravitational field of NSs unless the sexaquarks and other hadrons dissociate to quarks. The flavor content of sexaquarks $uuddss$ coincides with the one of three diquarks $ud$, $ds$ and $su$. Dissociation of sexaquarks causes formation of a color-superconducting three-flavor state of quark matter dominated by the above diquarks, i.e. the CFLL quark matter Blaschke et al. (2022). This state exists in thermodynamic equilibrium with the nuclear matter, i.e. their pressures and chemical potentials coincide. Thus, within the considered picture quark deconfinement can be modeled by matching quark and hadron EoSs by means of the Maxwell construction at $\mu_{B}=m_{S}/2$. This simultaneously respects the mentioned conditions of the BEC of sexaquarks and thermodynamic equilibrium.
@@
Providing the Maxwell matching of the EoSs of CFLL quark matter with the DD2npY-T one of hadron matter at $\mu_{B}=m_{S}/2$ requires adjusting the vector and diquark couplings of the nonlocal NJL model. Conventionally, below we treat $\eta_{V}$ as a single independent parameter of the model and adjust $\eta_{D}$ according to the above requirement. It is known that increasing the vector coupling increases the onset density of quark matter, while the diquark one causes the opposite effect (see the corresponding discussion e.g. in Refs. Baym et al. (2018); Ivanytskyi and Blaschke (2022a, b)). Therefore, having a fixed value of the chemical potential of quark deconfinement defined by the sexaquark mass requires $\eta_{D}$ to be a growing function of $\eta_{V}$. This behavior is seen from Table 1, which summarizes the hybrid EoSs used below. Since the considered scenario assumes deconfinement of a color-superconducting CFLL quark matter, the diquark coupling obtained from the requirement that quark-hadron transition onsets at the BEC of sexaquarks should provide a finite diquark pairing gap found as a solution of Eq. (25) at $\mu_{B}=m_{S}/2$. Within the considered scenario transition from hadron to quark matter in the general case is of the first order. It is indicated by a finite jump of the baryon density. Requiring this density jump across the transition from nuclear to quark matter to be non-negative, we limit the ranges of the vector and diquark couplings from above. The vector and diquark couplings producing the EoSs summarized in Table 1 respect these two requirements. They also assure the absence of the vacuum color superconductivity since they are always below the critical value $\eta_{D}=0.765$.
@@
Figure 2: Hybrid EoS of cold electrically neutral quark-hadron matter at $\beta$-equilibrium in the plane of pressure $p$ versus energy density $\varepsilon$. The curve corresponding to hadron matter is obtained without sexaquarks. The empty white circle indicates onset of hyperons in pure hadron matter. The calculations are performed for the sexaquark mass $m_{S}=2054$ MeV. The nuclear matter Krüger et al. (2013), astrophysical Miller et al. (2021); Hebeler et al. (2013) and perturbative QCD Fraga et al. (2014) constraints represented by the shaded areas are discussed in the text.
In order to model NSs we construct a family of hybrid quark-hadron EoSs based on the procedure described above. The resulting EoSs in the plane of pressure versus energy density are shown in Fig. 2. They agree well with the constraints extracted from the observational data of the pulsars PSR J1614+2230 Hebeler et al. (2013) and PSR J0740+6620 Miller et al. (2021). An insignificant tension with the PSR J0740+6620 constraint is due to the absence of a first order phase transition in the analysis of Ref. Miller et al. (2021), while the corresponding observational data are fully reproduced within the present model. The flat parts of the depicted curves correspond to the mixed quark-hadron phase where the BEC of sexaquarks exists. As is seen, increasing the vector coupling stiffens the quark branch of the EoS and simultaneously reduces the density jump across the quark-hadron transition. The maximal considered value of $\eta_{V}=1.21$ corresponds to vanishing of this jump. At sufficiently large values of this parameter the quark branch of the hybrid EoS is stiffer than the DD2npY-T EoS at the energy densities above $400~{}{\rm MeV~{}fm}^{-3}$. As is shown below, this allows the hybrid EoSs to reach the maximum NS masses above the value $2.1~{}{\rm M}_{\odot}$ corresponding to the DD2npY-T EoS. It is also seen from Fig. 2 that the developed EoSs are consistent with the constraint of perturbative QCD Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018); Fernandez and Kneur (2022). This is provided by the nonlocal treatment of quark interactions allowing the present model to asymptotically reach the conformal limit of QCD.
@@
The speed of sound and dimensionless interaction measure of the present model are shown in Fig. 3. The hadron branch of the hybrid EoS demonstrates monotonously growing $c_{S}^{2}$ and decreasing $\delta$. In the purely hadronic case $c_{S}^{2}$ is nonmonotonous due to the onset of hyperons. In the mixed phase region the speed of sound vanishes due to the constant pressure while the dimensionless interaction measure exhibits a non-smooth behavior caused by a discontinuous density jump. Note that at the limiting value of the vector coupling $\eta_{V}=1.21$ the first order quark-hadron transition degenerates to the second order one indicated by the vanishing density jump. In this case the region with vanishing $c_{S}^{2}$ disappears, while preserving a discontinuity of this quantity. At this value of the vector coupling $\delta$ is a smooth function of $\varepsilon$. The quark branch of the hybrid EoS exhibits a nonmonotonous behavior of the speed of sound and dimensionless interaction measure. Right after the quark deconfinement $c_{S}^{2}$ increases and $\delta$ decreases with growing energy density. This signals stiffening of the EoS caused by the quark deconfinement. The larger the vector coupling, the stiffer the quark branch of the EoS. This is manifested by increasing $c_{S}^{2}$ and decreasing $\delta$ at growing $\eta_{V}$. The speed of sound develops a peak at $\varepsilon\simeq 500~{}{\rm MeV/fm}^{3}$. Thus, $c_{S}^{2}$ reaches its maximal value in the quark matter phase,
@@
Precursors of such a behavior of $c_{S}^{2}$ and $\delta$ in the centers of the heaviest NSs were discussed within the context of approximately restored conformal symmetry Fujimoto et al. (2022); Annala et al. (2023). At the same time, the conformality of a theory assumes that the corresponding action is scale invariant. This implies an absence of dimensionful parameters in the action and leads to $p\propto\varepsilon\propto\mu^{4}$ due to the dimensionality reason. In this case $c_{S}^{2}=1/3$ and $\delta=0$ follow as a consequence of scale invariance. However, the inverse statement that $c_{S}^{2}=1/3$ and $\delta=0$ at a certain density imply the scale invariance is not correct in the general case. A simple counterexample is given by the model of ultraheavy quarks with local vector repulsion and a constant negative bag-like pressure shift. This model obviously breaks the scale invariance by three dimensionful parameters, i.e. by the quark mass, the vector coupling and the bag constant. The interested readers can find the corresponding consideration in Appendix A. At the same time, at the non-perturbative regimes the CFLL quark matter is characterized by the effective quark mass, vector field amplitude and pairing gap. These dynamically generated parameters are not negligible compared to the quark chemical potential and break the scale invariance of the effective action. For example, at $\eta_{V}=1$ and $\varepsilon=2100~{}\rm MeV/fm^{3}$ one gets $\omega/\mu=0.58$, while the speed of sound and dimensionless interaction measure are close to the conformal values. Therefore, the question about an approximately conformal behavior of quark matter at non-perturbative densities can not be addressed based just on the analysis of its EoS and requires an independent consideration of the effective action and relevant microscopic quantities. Such a study is beyond the scope of the present paper and should be performed separately.
@@
Fig. 3 also demonstrates that the scenario of the CFLL quark matter in NS agrees with the constraints on the speed of sound and dimensionless interaction measure at the energy densities reached in NSs. These constraints were extracted from the astrophysical data based on the model agnostic statistical analysis at the 94 % confidence level Altiparmak et al. ; Marczenko et al. . The mentioned agreement is rather remarkable since unlike the present work the analysis from Refs. Altiparmak et al. ; Marczenko et al. does not explicitly include the possibility of a first order phase transition. A possible phenomenological explanation may be related to an early quark deconfinement at an onset mass below the masses of the analyzed NSs. In this case all these astrophysical objects lay on the quark branch of the mass-radius relation and, thus, are weakly sensitive to the details of the deconfinement phase transition. It is also seen from Fig. 3 that the speed of sound of the nonlocal NJL model is consistent with the perturbative QCD predictions Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018); Fernandez and Kneur (2022) at the energy densities above $\rm 10^{5}~{}MeV/fm^{3}$. In the case of the dimensionless interaction measure the agreement requires even larger $\varepsilon$. As discussed in Section III, this is due to the impossibility to adjust the formfactor according to the perturbative QCD results on the behavior of $c_{S}^{2}$ and $\delta$ at asymptotically high densities. A better agreement can be reached by a piecewise $g_{\bf k}$ being Gaussian at small momenta and behaving as $g_{\bf k}\propto\alpha_{s}/{\bf k}^{2}$ at large momenta Hell et al. (2011); Kashiwa et al. (2011).
@@
As was shown in Section II, the effective quark mass, vector field amplitude and pairing gap are negligible at high densities. This insures the conformal behavior of the CFLL quark matter with the speed of sound and dimensionless interaction measure assymptoticlly converging to the conformal values $c_{S}^{2}=1/3$ and $\delta=0$. Fig. 3 demonstrates this. It is also seen that convergence to the conformal limit is delayed by increasing the vector coupling. For the Gaussian formfactor and a finite current quark mass $c_{S}^{2}\simeq 1/3-m^{2}/3\mu^{2}$ and $\delta\simeq m^{2}/3\mu^{2}$ at high densities, while the terms corresponding to the vector repulsion are subdominant corrections. It follows from Eq. (35) that they can be neglected if $\mu^{3}g_{\mu}\ll-m^{2}\Lambda^{2}/\omega_{\infty}$. This condition is strongly dependent on the vector coupling appearing through the asymptotic value of the vector field. For example, at $\eta_{V}=0.55$ and $\eta_{V}=1.00$ this requires $\mu_{B}\gg 6439$ MeV, $\varepsilon\gg 5.10\cdot 10^{7}~{}\rm MeV/fm^{3}$ and $\mu_{B}\gg 6586$ MeV, $\varepsilon\gg 5.58\cdot 10^{7}~{}{\rm MeV/fm}^{3}$, respectively. At smaller chemical potentials and energy densities vector repulsion significantly affects the behavior of speed of sound and dimensionless interaction measure.
@@
Comparing the speed of sound and dimensionless interaction measure of the present model and the NJL model with local vector repulsion from Ref. Blaschke et al. (2022) allows us to analyze the phenomenological influence of the nonlocal treatment of the mentioned interaction channel. Fig. 3 shows $c_{S}^{2}$ and $\delta$ of pure quark matter obtained for the model with the local vector interaction. The first order phase transition from chirally broken normal quark matter to partially chirally symmetric color-superconducting quark matter is indicated in in Fig. 3 by discontinuous jumps of $c_{S}^{2}$ and kinks of $\delta$. In the case of local treatment of the vector repulsion the NJL model produces rather stiff EoS of quark matter. This leads to large values of the speed of sound always exceeding the one corresponding to the nonlocal NJL model. As is seen, unlike the nonlocal NJL model, $c_{S}^{2}$ monotonously grows asymptotically reaching the speed of light (see the discussion in Section III). As argued above, the local treatment of the vector repulsion does not allow the NJL model to reach the conformal limit. The same conclusion can be drawn from the behavior of the dimensionless interaction measure. At the local vector repulsion it monotonously decreases saturating to the nonconformal value $-2/3$. It is also worth mentioning that in the case of the local treatment of the vector repulsion the quark matter EoS does not exhibit a region where $c_{S}^{2}$ and $\delta$ simultaneously attain nearly conformal values.